# A Dynamical View of the Question of Why

### 14 Feb 2024 v1

@@ 5 Experiments
In this experiment, we analyze multivariate medical time-series data using an open-source implementation of the FDA-approved Type-1 Diabetes Mellitus Simulator (T1DMS) (Kovatchev et al., 2009). The simulator models a patient’s blood glucose (BG) level and 12 other real-valued signals representing glucose and insulin values in different body compartments. We control two actions: 1) insulin intake, to regulate insulin level 2) meal intake, to manage the amount of carbohydrates. More details on the data and experiment design can be found in Appendix E. This real-world simulation helps us investigate the impact of insulin and carbohydrate intake on blood glucose levels. Meal intake increases the amount of glucose in the bloodstream, which for T1D patients, is regulated with external insulin intake. However, intensive control of blood glucose with insulin injections can increase the risk of hypoglycemia (BG level < 70 mg/dL), which is the effect event we aim to understand its causes (event $B$). We use Monte Carlo learning to estimate $V_{\Gamma}(\textbf{X})$ and thus $\Gamma_{B}(\textbf{X})$. Of note, the algorithm has only access to data. In our setting, we observe that intake of insulin causes an instantaneous spike in dynamics of subcutaneous insulin 1 (SI1), while intake of carbohydrates causes an instantaneous spike in glucose in stomach 1 (GS1). Since the previous action is considered as part of the current state, the spike in subcutaneous insulin just acts as a proxy for action insulin, similarly for action meal and GS1. We study two scenarios under event $B$:
@@ Acknowledgments
We express our sincere gratitude to our colleagues whose invaluable advice enhanced the quality of this work. Special thanks are extended to the (last) RL team and other colleagues at MSR Montreal for their pivotal suggestions during the early stages of this project. We are particularly grateful for the insightful feedback offered by Marzyeh Ghassemi, Elliot Creager, and Rahul G. Krishnan. Additionally, we would like to acknowledge the anonymous reviewers for their encouraging remarks and constructive suggestions, which helped to improve this paper.
@@ C.1 Overview
  • Time is an explicit factor in our formulation. It establishes the direction of causation between events and answer questions of causation in more practical and complicated scenarios. Since we define events as changes in variables over a homogeneous interval of time, it clears much of the confusion around the time of happening of an event and provides the flexibility to study multiple events that share the same ruling variables and admit the same changes but occur at different points in time.
@@ C.2 Structural Equation Modelling
This section reviews the definition of causal models as they were introduced by Pearl in [Pearl et al., 2000]. We provide this here so we are equipped with the basic tools needed to understand the HP definition. Much of the discussion and notation is taken from Halpern’s actual causality [Halpern, 2016], as presented in Beckers [2021] with little change.
@@
A signature $\mathcal{P}$ is a tuple $(\mathcal{U},\mathcal{V},\mathcal{R})$, where $\mathcal{U}$ is a set of exogenous variables, $\mathcal{V}$ is a set of endogenous variables and $\mathcal{R}$ a function that associates with every variable $Y\in\mathcal{U}\cup\mathcal{V}$ a nonempty set $\mathcal{R}(Y)$ of possible values for $Y$ (i.e., the set of values over which $Y$ ranges). If $\vec{X}=\left(X_{1},\ldots,X_{n}\right),\mathcal{R}(\vec{X})$ denotes the cross product $\mathcal{R}\left(X_{1}\right)\times\cdots\times\mathcal{R}\left(X_{n}\right)$.
Exogenous variables represent factors whose causal origins are outside the scope of the causal model, such as background conditions and noise. The values of the endogenous variables, on the other hand, are causally determined by other variables within the model (both endogenous and exogenous).
@@
A causal model $M$ is a pair $(\mathcal{P},\mathcal{F})$, where $\mathcal{P}$ is a signature and $\mathcal{F}$ defines a function that associates with each endogenous variable $X$ a structural equation $F_{X}$ giving the value of $X$ in terms of the values of other endogenous and exogenous variables. Formally, the equation $F_{X}$ maps $\mathcal{R}(\mathcal{U}\cup\mathcal{V}-\{X\})$ to $\mathcal{R}(X)$, so $F_{X}$ determines the value of $X$, given the values of all the other variables in $\mathcal{U}\cup\mathcal{V}$.
Note that there are no functions associated with exogenous variables; their values are determined outside the model. We call a setting $\vec{u}\in\mathcal{R}(\mathcal{U})$ of values of exogenous variables a context.
The value of $X$ may depend on the values of only a few other variables. $X$ depends on $Y$ in context $\vec{u}$ if there is some setting of the endogenous variables other than $X$ and $Y$ such that if the exogenous variables have value $\vec{u}$, then varying the value of $Y$ in that context results in a variation in the value of $X$; that is, there is a setting $\vec{z}$ of the endogenous variables other than $X$ and $Y$ and values $y$ and $y^{\prime}$ of $Y$ such that $F_{X}(y,\vec{z},\vec{u})\neq F_{X}\left(y^{\prime},\vec{z},\vec{u}\right)$. We then say that $Y$ is a parent of $X$. A causal path is a sequence of variables in which each element is a child of the previous element.
A causal model is strongly recursive (or strongly acyclic) if there is a partial order $\prec$ on variables such that if Y depends on X, then $X\preceq Y$. In a strongly recursive model, given a context $\vec{u}$, the values of all the remaining variables are determined (we can just solve for the value of the variables in the order given by $\preceq$).
An intervention has the form $\vec{X}\leftarrow\vec{x}$, where $\vec{X}$ is a set of endogenous variables. Intuitively, this means that the values of the variables in $\vec{X}$ are set to the values $\vec{x}$. The structural equations define what happens in the presence of interventions. Setting the value of some variables $\vec{X}$ to $\vec{x}$ in a causal model $M=(\mathcal{P},\mathcal{F})$ results in a new causal model, denoted $M_{\vec{X}\leftarrow\vec{x}}$, which is identical to $M$, except that $\mathcal{F}$ is replaced by $\mathcal{F}^{\vec{X}\leftarrow\vec{x}}$ : for each variable $Y\notin\vec{X},F_{Y}^{\vec{X}\leftarrow\vec{x}}=F_{Y}$ (i.e., the equation for $Y$ is unchanged), while for each $X^{\prime}$ in $\vec{X}$, the equation $F_{X^{\prime}}$ for $X^{\prime}$ is replaced by $X^{\prime}=x^{\prime}$ (where $x^{\prime}$ is the value in $\vec{x}$ corresponding to $\left.X^{\prime}\right)$.
Given a signature $\mathcal{P}=(\mathcal{U},\mathcal{V},\mathcal{R})$, an atomic formula is a formula of the form $X=x$, for $X\in\mathcal{V}$ and $x\in\mathcal{R}(X)$. A causal formula (over $\mathcal{P}$ ) is one of the form $\left[Y_{1}\leftarrow y_{1},\ldots,Y_{k}\leftarrow y_{k}\right]\varphi$, where
  • $\varphi$ is a Boolean combination of atomic formulas,
  • $Y_{1},\ldots,Y_{k}$ are distinct variables in $\mathcal{V}$, and
  • $y_{i}\in\mathcal{R}\left(Y_{i}\right)$ for each $1\leq i\leq k$.
Such a formula is abbreviated as $[\vec{Y}\leftarrow\vec{y}]\varphi$. The special case where $k=0$ is abbreviated as $\varphi$. Intuitively, $\left[Y_{1}\leftarrow y_{1},\ldots,Y_{k}\leftarrow y_{k}\right]\varphi$ says that $\varphi$ would hold if $Y_{i}$ were set to $y_{i}$, for $i=1,\ldots,k$.
A causal formula $\psi$ is true or false in a causal setting, which is a causal model given a context. As usual, we write $(M,\vec{u})\models\psi$ if the causal formula $\psi$ is true in the causal setting $(M,\vec{u})$. The relation is defined inductively. $(M,\vec{u})\models X=x$ if the variable $X$ has value $x$ in the unique (since we are dealing with recursive models) solution to the equations in $M$ in context $\vec{u}$ (i.e., the unique vector of values that simultaneously satisfies all equations in $M$ with the variables in $\mathcal{U}$ set to $\vec{u}$ ).
@@ C.3 HP Definitions
Given we are equipped with the background of structural equation modeling, we can now move on to the HP definitions. As Pearl’s [Pearl et al., 2000] initial definition, is a precursor to the HP definitions, gives less intuitive results, and is far more complicated, we do not discuss it. (It is safe to say that by now it has been unanimously rejected.) Two of the HP definitions are developed by both Halpern and Pearl [Halpern and Pearl, 2001, 2005], whereas the third one is solely due to Halpern [Halpern, 2015]. The relations between them are extensively discussed by Halpern Halpern [2016]. This discussion is heavily borrowed from Halpern [2016] and Beckers [2021].
Questions of actual causation are posed relative to an actual context $\vec{u}$, which determines which events actually took place. Throughout our discussion, settings of variables $V$ with superscript *** (i.e., $\left(\vec{v}^{*}\right)$ indicate that $(M,\vec{u})\models\left(\vec{V}=\vec{v}^{*}\right)$. Settings of variables $\vec{V}$ with superscript ${}^{\prime}$ (i.e., $\left.\vec{v}^{\prime}\right)$ indicate that $(M,\vec{u})\models\left(V\neq v^{\prime}\right)$ for each $V\in\vec{V}$. Settings of variables without any superscript can refer to any setting.
Given this understanding, the original HP definition [Halpern and Pearl, 2001] are as follows:
Definition 4 (Original HP [Halpern and Pearl, 2001]).
@@
    There is a partition of $\mathcal{V}$ into two sets $\vec{Z}$ and $\vec{W}$ with $\vec{X}\subseteq\vec{Z}$ and a setting $\vec{x}^{\prime}$ and $\vec{w}$ of the variables in $\vec{X}$ and $\vec{W}$, respectively, such that $(M,\vec{u})\models\left[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}\right]\neg\varphi$
@@
    For all subsets $\vec{O}$ of $\vec{Z}-\vec{X}$, we have $(M,\vec{u})\models\left[\vec{X}\leftarrow\vec{x},\vec{W}\leftarrow\vec{w},\vec{O}\leftarrow\vec{o}^{*}\right]\varphi$.
@@
    $\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime\prime}=\vec{x}^{\prime\prime}$ satisfies AC2, where $\vec{x}^{\prime\prime}$ is the restriction of $\vec{x}$ to the variables in $\vec{X}^{\prime\prime}$.
It’s important to note that in the above definition, we call $\vec{W}=\vec{w}$ a witness of $\vec{X}=\vec{x}$ causing $\varphi$. Variables in $\vec{Z}$ are thought of as making up the “causal path” from $\vec{X}$ to $\varphi$. Intuitively, changing the value of some variable in $\vec{X}$ results in changing the value(s) of some variable(s) in $\vec{Z}$, which results in the values of some other variable(s) in $\vec{Z}$ being changed, which finally results in the truth value of $\varphi$ changing.
AC1 represents the trivial requirement that the candidate cause and effect are among the events which took place. AC3 is also fairly straightforward: we should not consider redundant elements to be parts of causes.
The real content of the definition lies with AC2. Halpern refers to conditions AC2(a) and AC2(b) as "necessity condition" and "sufficiency condition" [Halpern, 2015]. AC2(a) states that the effect counterfactually depends on the cause when holding fixed the witness $\vec{W}=\vec{w}:(M,\vec{u})\models[\vec{X}\leftarrow$ $\vec{x},\vec{W}\leftarrow\vec{w}]\varphi$ and $(M,\vec{u})\models\left[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}\right]\neg\varphi$. Therefore AC2(a) can easily be interpreted as expressing a - contrastive - necessity condition: there exist contrast values $\vec{x}^{\prime}$ such that if $\vec{X}$ is set to $\vec{x}^{\prime}$, $\varphi$ no longer holds.
The sufficiency condition, AC2(b), roughly speaking, requires that if the variables in $\vec{X}$ and an arbitrary subset $\vec{O}$ of other variables on the causal path (i.e., besides those in $\vec{X}$ ) are held at their values in the actual context (where the values of the variables in $\vec{X}$ in the actual context is $\vec{x}$, by AC1), then $\varphi$ holds even if $\vec{W}$ is set to $\vec{w}$ (the setting for $\vec{W}$ used in AC2(a)). Note that, due to setting $\vec{W}$ to $\vec{w}$, the values of the variables in $\vec{Z}$ may change. AC2(b) says that this change does not affect $\varphi$; $\varphi$ continues to be true.
The problem lies with interpreting AC2(b) as expressing causal sufficiency. The main obstacle lies in the absence of the requirement that $\vec{w}=\vec{w}^{*}$, i.e., it is not required that the supposedly sufficient set of events $(\vec{X}=\vec{x},\vec{W}=\vec{w})$ actually took place.
Definition 5 (Updated HP [Halpern and Pearl, 2005]).
Keeping AC1 and AC3 the same, condition AC2 is stated as follows:
  • AC2(a).
    Identical to original HP.
  • AC2(b).
    For all subsets $\vec{Q}$ of $\vec{W}$ and subsets $\vec{O}$ of $\vec{Z}-\vec{X}$, we have $(M,\vec{u})\models\big{[}\vec{X}\leftarrow$ $\vec{x},\vec{Q}\leftarrow\vec{q},\vec{O}\leftarrow\vec{o}^{*}\big{]}\varphi~{}~{}$.
The only difference between AC2(b) in the updated and the original version lies in the clause "for all subsets $\vec{Q}$ of $\vec{W}$": AC2(b) in the updated version must hold even if only a subset $\vec{Q}$ of the variables in $\vec{W}$ are set to their values in $\vec{w}$. Given modified condition AC2(a), this means that the variables in $\vec{W}-\vec{Q}$ let’s say $\vec{W^{\prime}}$, essentially act as they do in the real world; that is, their values are determined by the structural equations, $*\vec{w^{{}^{\prime}*}}$. Halpern says this definition is motivated by the intuition that only what happens in the actual situation should matter.
Definition 6 (Modified HP [Halpern, 2015]).
Keeping conditions AC1 and AC3 the same, condition AC2 is stated as:
  • AC2(a).
    There is a set $\vec{W}$ of variables in $\mathcal{V}-\vec{X}$, and a setting $\vec{x}^{\prime}$ of the variables in $\vec{X}$ such that $(M,\vec{u})\models\big{[}\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}^{*}\big{]}\neg\varphi$.
  • AC2(b).
    Identical to updated HP.
Here counterfactual dependence of $\phi$ on $\vec{X}$ is explored by keeping the variables in $\vec{W}$ fixed at their actual values. The crucial difference here is that Modified HP does require the witness to consist solely of events that actually took place, i.e., $\vec{w}=\vec{w}^{*}$. It is straightforward to see that simply by adding this requirement both versions of AC2(b) are satisfied automatically, and therefore an explicit sufficiency condition is not required. Halpern considers this definition to be an improvement over the other two and mentions that the need for a sufficiency condition arises only if we are considering contingencies that differ from the actual setting in AC2(a).
@@ C.5 Compare and Contrast with HP definition
Given the current understanding of the HP definitions, let us now compare and contrast the modified HP definition of actual causation (Def. 6) with our definition of causation (Def.1). For the purposes of this discussion, let’s consider $\vec{X}=\vec{x}$ as event $A$ and $\varphi$ or $\vec{Y}=\vec{y}$ as event $B$. Also, since the HP definitions do not explicitly consider time in the definition of events, for simplicity let’s assume that variables $\vec{X}$ admit values $\vec{x}$ only once during the observation period. The same applies to $\vec{Y}=\vec{y}$. From discussed mapping in C.4, $\vec{X}=\vec{x}$ implies $S_{\mathcal{D}_{A}}$, admits a certain change of values between $[t_{A},t^{\prime}_{A}]$, say $S_{\mathcal{D}_{A}}(t^{\prime}_{A})=\vec{x}$ (implying that in this case, for simplicity, we consider that $S_{\mathcal{D}_{A}}(t_{A})$ admits some value of no consequential interest here). Similar logic can be applied to event $B$.
HP Definition:
$\vec{X}=\vec{x}$ is an actual cause of $\varphi$, $\vec{Y}=\vec{y}$ in $(M,\vec{u})$ if the following three conditions hold:
  1. AC1.
    $(M,\vec{u})\models(\vec{X}=\vec{x})\wedge\varphi$
  2. AC2(a).
  3. AC2(b).
    For all subsets $\vec{Q}$ of $\vec{W}$ and subsets $\vec{O}$ of $\vec{Z}-\vec{X}$, we have $(M,\vec{u})\models\big{[}\vec{X}\leftarrow$ $\vec{x},\vec{Q}\leftarrow\vec{q}^{*},\vec{O}\leftarrow\vec{o}^{*}\big{]}\varphi~{}~{}$ (where $\vec{q}$ is the restriction of $\vec{w}$ to $\vec{Q}$).
  4. AC3.
    $\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime\prime}=\vec{x}^{\prime\prime}$ satisfies AC2, where $\vec{x}^{\prime\prime}$ is the restriction of $\vec{x}$ to the variables in $\vec{X}^{\prime\prime}$.
@@
AC1 represents the trivial requirement that the candidate cause and effect are among the events that took place. Our condition C1. implicity covers HP condition AC1 (Def. 6), adding additional clarity of direction of causation with a time arrow between the events.
@@
Comparison: We argue that the intent of condition AC2(a) in Def.6 is to understand the effect of event $A$ ($\vec{X}=\vec{x}$ or $S_{\mathcal{D}_{A}}(t_{A})=\vec{x}$) in isolation. This condition looks at the effect of the non-happening of event $A$ ($\vec{X}=\vec{x}^{\prime}$ or $S_{\mathcal{D}_{A}}(t_{A})=\vec{x}^{\prime}$) on event $B$, while witness $\vec{W}$ are held under the values of context $\vec{u}$, i.e, when event $B$ ($\vec{Y}=\vec{y}$) occurs. Our conditions C2 and C3 address this issue without having to take an interventionist approach to causation. In our formulation, we can compute the contribution of each event (defined by a subset of the state/action components over a homogenous time interval) towards increasing the minimum probability of the event $B$ happening. This helps us talk about the individual contribution of each event of interest without the need to hold $S_{\mathcal{D}_{W}}(t_{W})$ at the value achieved under context $\vec{u}$ while admitting a different value at $S_{\mathcal{D}_{A}}(t_{A})$, which might be physically impossible to achieve.
@@
The sufficiency condition, as discussed in condition AC2(b) of Modified HP (Def. 6) roughly speaking mentions that if the variables in $\vec{X}$ and an arbitrary subset $\vec{O}$ of other variables on the causal path are held at their values in the actual context $\vec{u})$ (i.e, $\vec{X}(u)=\vec{x}$ and $\vec{O}(u)=\vec{o}^{*}$), then $\varphi$ holds even if any subset of $\vec{W}$, $\vec{Q}$ is set to $\vec{q}^{*}$. Therefore, this condition implies that once event $\vec{X}=\vec{x}$ happens, then no matter what values the events on the causal path between event $\vec{X}=\vec{x}$ and event $\varphi$, ($\vec{Y}=\vec{y}$) take, event $\vec{Y}=\vec{y}$ still happens.

### 27 Feb 2024 v2

@@ 5 Experiments
In this experiment, we analyze multivariate medical time-series data using an open-source implementation of the FDA-approved Type-1 Diabetes Mellitus Simulator (T1DMS) (Kovatchev et al., 2009). The simulator models a patient’s blood glucose (BG) level and 12 other real-valued signals representing glucose and insulin values in different body compartments. We control two actions: 1) insulin intake, to regulate insulin level 2) meal intake, to manage the amount of carbohydrates. More details on the data and experiment design can be found in Appendix E. This experiment helps us investigate the impact of insulin and carbohydrate intake on blood glucose levels. Meal intake increases the amount of glucose in the bloodstream, which for T1D patients, is regulated with external insulin intake. However, intensive control of blood glucose with insulin injections can increase the risk of hypoglycemia (BG level < 70 mg/dL), which is the effect event we aim to understand its causes (event $B$). We use Monte Carlo learning to estimate $V_{\Gamma}(\textbf{X})$ and thus $\Gamma_{B}(\textbf{X})$. Of note, the algorithm has only access to data. In our setting, we observe that intake of insulin causes an instantaneous spike in dynamics of subcutaneous insulin 1 (SI1), while intake of carbohydrates causes an instantaneous spike in glucose in stomach 1 (GS1). Since the previous action is considered as part of the current state, the spike in subcutaneous insulin just acts as a proxy for action insulin, similarly for action meal and GS1. We study two scenarios under event $B$:
@@ Acknowledgments
We express our sincere gratitude to our colleagues whose invaluable advice enhanced the quality of this work. Special thanks are extended to the former RL team and other colleagues at MSR Montreal for their pivotal suggestions during the early stages of this project. We are particularly grateful for the insightful feedback offered by Marzyeh Ghassemi, Elliot Creager, and Rahul G. Krishnan. Additionally, we would like to acknowledge the anonymous reviewers for their encouraging remarks and constructive suggestions, which helped to improve this paper.
@@ C.1 Overview
  • Time is an explicit factor in our formulation. It establishes the direction of causation between events and answers questions of causation in more practical and complicated scenarios. Since we define events as changes in variables over a homogeneous interval of time, it clears much of the confusion around the time of happening of an event and provides the flexibility to study multiple events that share the same ruling variables and admit the same changes but occur at different points in time.
@@ C.2 Structural Equation Modelling
Before we look into the HP definitions we assume that the reader is familiar with the concept of structural equation modeling. To do so, we recommend the reader refer to [Beckers, 2021] section 2, so they are equipped with the basic tools needed to understand the HP definition. For our discussion, we will borrow heavily from this reference for notions and discussions necessary to define and understand structural causal modeling.
Much of the discussion and notation is taken from Halpern’s actual causality [Halpern, 2016], as presented in Beckers [2021] with little change.
@@
[Beckers, 2021] A signature $\mathcal{P}$ is a tuple $(\mathcal{U},\mathcal{V},\mathcal{R})$, where $\mathcal{U}$ is a set of exogenous variables, $\mathcal{V}$ is a set of endogenous variables and $\mathcal{R}$ a function that associates with every variable $Y\in\mathcal{U}\cup\mathcal{V}$ a nonempty set $\mathcal{R}(Y)$ of possible values for $Y$ (i.e., the set of values over which $Y$ ranges). If $\vec{X}=\left(X_{1},\ldots,X_{n}\right),\mathcal{R}(\vec{X})$ denotes the cross product $\mathcal{R}\left(X_{1}\right)\times\cdots\times\mathcal{R}\left(X_{n}\right)$.
It’s important to recognize that exogenous variables are factors whose causes lie beyond the direct influence of the model, encompassing background conditions and noise. Conversely, the values of endogenous variables are causally influenced by other variables within the model, whether they are endogenous or exogenous. We refer to a collection $\vec{u}\in\mathcal{R}(\mathcal{U})$ of exogenous variable values as the contextual setup of the problem.
@@
[Beckers, 2021] A causal model $M$ is a pair $(\mathcal{P},\mathcal{F})$, where $\mathcal{P}$ is a signature and $\mathcal{F}$ defines a function that associates with each endogenous variable $X$ a structural equation $F_{X}$ giving the value of $X$ in terms of the values of other endogenous and exogenous variables. Formally, the equation $F_{X}$ maps $\mathcal{R}(\mathcal{U}\cup\mathcal{V}-\{X\})$ to $\mathcal{R}(X)$, so $F_{X}$ determines the value of $X$, given the values of all the other variables in $\mathcal{U}\cup\mathcal{V}$.
@@ C.3 HP Definitions
We now look at HP definitions of causation as presented in [Halpern, 2015]. Halpern and Pearl [Halpern and Pearl, 2001, 2005] develop two of the initial HP definitions, whereas the third one is proposed solely by Halpern [Halpern, 2015]. This will be the one we will discuss in detail since it builds on the limitations of the earlier proposed definitions. The relations between them are extensively discussed by Halpern Halpern [2016]. We encourage the readers to read through this work to get a better insight into it.
We borrow on the notions used by [Beckers, 2021] in section 3 to discuss the HP definitions. Throughout our discussion, settings of variables $V$ with
  • *** i.e., $\vec{v}^{*}$ indicate that $(M,\vec{u})\models\left(\vec{V}=\vec{v}^{*}\right)$.
  • ${}^{\prime}$ i.e., $\vec{v}^{\prime}$ indicate that $(M,\vec{u})\models\left(\vec{V}\neq\vec{v}^{\prime}\right)$.
  • No subscripts can refer to any setting.
Given the notations, the modified HP definition Halpern [2015] is given as follows:
Definition 4 (Modified HP [Halpern, 2015]).
@@
    For all subsets $\vec{Q}$ of $\vec{W}$ and subsets $\vec{O}$ of $\vec{Z}-\vec{X}$, we have $(M,\vec{u})\models\big{[}\vec{X}\leftarrow$ $\vec{x},\vec{Q}\leftarrow\vec{q},\vec{O}\leftarrow\vec{o}^{*}\big{]}\varphi~{}~{}$
@@
    $\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime\prime}=\vec{x}^{\prime\prime}$ satisfies AC2, where $\vec{x}^{\prime\prime}$ is the restriction of $\vec{x}$ to the variables in $\vec{X}^{\prime\prime}$
We designate $\vec{W}=\vec{w}$ as a witness indicating that when $\vec{X}=\vec{x}$, it causes the occurrence of $\varphi$. The variables within $\vec{Z}$ are conceptualized as constituting the "causal path" from $\vec{X}$ to $\varphi$. In essence, altering the value of a variable in $\vec{X}$ leads to changes in the value(s) of certain variable(s) in $\vec{Z}$, subsequently influencing the values of other variable(s) within $\vec{Z}$, ultimately resulting in a change in the truth value of $\varphi$.
AC1 stipulates the basic criterion that both the potential cause and effect must be events that occurred. AC3 is similarly straightforward, emphasizing the exclusion of redundant elements from the causal factors. However, the crux of the definition lies in AC2. Halpern categorizes conditions AC2(a) and AC2(b) as the "necessity condition" and the "sufficiency condition," respectively [Halpern, 2015].
AC2(a) asserts that the effect demonstrates counterfactual dependence on the cause, holding the witness fixed at their actual values, i.e, $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x},\vec{W}\leftarrow\vec{w}^{*}]\varphi$, and $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}^{*}]\neg\varphi$. Consequently, AC2(a) can be interpreted as articulating a contrastive necessity condition: there exist contrasting values $\vec{x}^{\prime}$ such that altering $\vec{X}$ to $\vec{x}^{\prime}$ results in non-fulfillment of $\varphi$.
The sufficiency condition, AC2(b), essentially stipulates that when the variables in $\vec{X}$ and any chosen subset $\vec{O}$ of other variables along the causal path (besides those in $\vec{X}$) retain their actual context values, then $\varphi$ holds even if only a subset $\vec{Q}$ of the variables in $\vec{W}$ are set to their values in $\vec{w}$ (the setting for $\vec{W}$ utilized in AC2(a)). Notably, by fixing $\vec{W}$ to $\vec{w}$, alterations in the values of variables within $\vec{Z}$ may occur. AC2(b) asserts that these changes do not impact the truth of $\varphi$; it remains valid. Consequently, in light of condition AC2(a), this implies that the variables in $\vec{W}-\vec{Q}$, denoted as $\vec{W^{\prime}}$, essentially operate as they do in reality; their values are determined by the structural equations, denoted as $*\vec{w^{{}^{\prime}*}}$.
@@ C.5 Compare and Contrast with HP definition
Given the current understanding of the HP definitions, let us now compare and contrast the modified HP definition of actual causation (Def. 4) with our definition of causation (Def.1). For the purposes of this discussion, let’s consider $\vec{X}=\vec{x}$ as event $A$ and $\varphi$ or $\vec{Y}=\vec{y}$ as event $B$. Also, since the HP definitions do not explicitly consider time in the definition of events, for simplicity let’s assume that variables $\vec{X}$ admit values $\vec{x}$ only once during the observation period. The same applies to $\vec{Y}=\vec{y}$. From discussed mapping in C.4, $\vec{X}=\vec{x}$ implies $S_{\mathcal{D}_{A}}$, admits a certain change of values between $[t_{A},t^{\prime}_{A}]$, say $S_{\mathcal{D}_{A}}(t^{\prime}_{A})=\vec{x}$ (implying that in this case, for simplicity, we consider that $S_{\mathcal{D}_{A}}(t_{A})$ admits some value of no consequential interest here). Similar logic can be applied to event $B$.
@@
AC1 represents the trivial requirement that the candidate cause and effect are among the events that took place. Our condition C1. implicity covers HP condition AC1 (Def. 4), adding additional clarity of direction of causation with a time arrow between the events.
@@
Comparison: We argue that the intent of condition AC2(a) in Def.4 is to understand the effect of event $A$ ($\vec{X}=\vec{x}$ or $S_{\mathcal{D}_{A}}(t_{A})=\vec{x}$) in isolation. This condition looks at the effect of the non-happening of event $A$ ($\vec{X}=\vec{x}^{\prime}$ or $S_{\mathcal{D}_{A}}(t_{A})=\vec{x}^{\prime}$) on event $B$, while witness $\vec{W}$ are held under the values of context $\vec{u}$, i.e, when event $B$ ($\vec{Y}=\vec{y}$) occurs. Our conditions C2 and C3 address this issue without having to take an interventionist approach to causation. In our formulation, we can compute the contribution of each event (defined by a subset of the state/action components over a homogenous time interval) towards increasing the minimum probability of the event $B$ happening. This helps us talk about the individual contribution of each event of interest without the need to hold $S_{\mathcal{D}_{W}}(t_{W})$ at the value achieved under context $\vec{u}$ while admitting a different value at $S_{\mathcal{D}_{A}}(t_{A})$, which might be physically impossible to achieve.
@@
The sufficiency condition, as discussed in condition AC2(b) of Modified HP (Def. 4) roughly speaking mentions that if the variables in $\vec{X}$ and an arbitrary subset $\vec{O}$ of other variables on the causal path are held at their values in the actual context $\vec{u})$ (i.e, $\vec{X}(u)=\vec{x}$ and $\vec{O}(u)=\vec{o}^{*}$), then $\varphi$ holds even if any subset of $\vec{W}$, $\vec{Q}$ is set to $\vec{q}^{*}$. Therefore, this condition implies that once event $\vec{X}=\vec{x}$ happens, then no matter what values the events on the causal path between event $\vec{X}=\vec{x}$ and event $\varphi$, ($\vec{Y}=\vec{y}$) take, event $\vec{Y}=\vec{y}$ still happens.